Question #3 You want to rent an unfurnished one-bedroom apartment in Boston next year. A 95% confidence interval for the true mean monthly rent of all one-bedroom apartments in Boston is (1263.64, 1536.36). What is the margin of error for this interval?
Question #4 A student curious about the average number of chocolate chips in a commercial brand of cookie estimated the standard deviation to be 8. If he wants to be 99% confident in his results, how manychocolate chip cookies will he need to sample to estimate the mean to within 2?
1
Expert's answer
2019-11-15T09:59:05-0500

1.


"margin\ of\ error={1 \over 2}\times (length\ of\ the\ interval)"

"margin\ of\ error={1 \over 2}\times(1536.36-1263.64)"

"margin\ of\ error=136.36"

2.Based on the information provided, the significance level is "\alpha=0.01" ("99\%" confidence interval), and the critical value for a two-tailed test is "z_c=2.58"


"margin\ of\ error=z_c\times{\sigma \over\sqrt{n}}"

Given that "\sigma=8, margin\ of\ error\leq2." Then


"2.58\times{8 \over\sqrt{n}}\leq2"

"n\geq({2.58\times8 \over 2})^2"

"n\geq106.5"

Since "n" is integer, then "n\geq107."